# I’m Sorry Dave: How the old world of personnel security can inform the new world of AI insider risk

### v1

@@ Abstract
A human insider is said to be intentional (as opposed to unintentional or unwitting) if they deliberately perform illicit and potentially harmful actions despite knowing these actions to be illicit and potentially harmful. Mapping intentionality onto an LLM would mean equating the system’s intent with its design features or its instructions (prompts). LLM intentionality could also be an emergent behaviour which is not contained in the system’s design or explicit instructions. Intentionality might also be implicit in the system’s instructions to achieve a particular goal. For example, an LLM might be directed (through its design or prompts) to act deceptively; or it might be implicitly required to act deceptively in order to achieve certain goals; or it might ‘pivot’ to act deceptively in ways not foreseen or instructed by its designers or users. The table below illustrates some of the ways in which an LLM might be said to act intentionally or unintentionally, depending on the actions of the model’s human developers or users.
Unintentional AI behaviour Intentional AI behaviour
Developers 1) Developers do not predict how their AI systems could be exploited e.g. a system designed to generate convincing legal documents could be misused for fraud.
2) Developers do not consider how an AI system could use deceit to achieve its goals. For example, systems taught to play games based on human data may adopt strategies that appear deceptive. Systems may appear more deceitful over time, due to changing goals or data drift.
3) AI systems are developed that hide their true ambitions as a result of adversarial training data or as emergent properties of instruction tuning.
1) Developers intentionally optimise an AI system for persuasion or deception to meet its goals e.g. in marketing, propaganda, or fraud.
2) Developers design an AI system to be deceitful in detecting and targeting vulnerable users.
3) A contractor or third-party developer intentionally embeds a backdoor in an AI system, allowing its behaviour or goals to be covertly altered later.
4) Developers use obfuscation to hide the true goals and intentions of AI systems.
Users 1) Employees use an AI system to generate marketing communications without realising that their contents are inaccurate. 1) Disgruntled employees intentionally use an AI system to create fake customer complaints.
2) An insider subverts an AI system by realigning its goals against the organisation.
Table 1: How AI systems can be intentionally and unintentionally misused by developers and users.
@@
Human insiders vary in the extent to which their illicit and potentially harmful actions are self-starting and self-directed, as opposed to manipulated, directed, or coerced by an external threat actor such as a hostile foreign state or a malign human insider.
AI systems are also potentially vulnerable to being manipulated or directed by an external threat actor such as a hostile foreign state. [35]
@@
Human insiders are usually difficult to detect. Capable and intentional insiders act covertly because they do not wish to be caught. The best ones may never be found. The archetypal example is the spy working within a sensitive government organisation who secretly acts for years on behalf of a hostile foreign intelligence agency. One consequence of covertness is that the true extent of insider activity tends to be systematically under-estimated, as organisations mistake the absence of evidence for evidence of absence. An AI insider might be even more difficult to detect, for a variety of reasons. It will be faster than humans and better at ingesting and analysing huge quantities of information. Moreover, its training data is likely to have included every known insider case in history. Furthermore, an AI insider’s strangeness, from the perspective of humans, may add to its covertness by making it harder to understand.
AI systems could, in theory, also help human insiders to avoid detection. They could be subverted to analyse patterns of activity within an organisation and subsequently advise external threat actors (their handlers) how to conduct effective social engineering attacks aimed at exfiltrating information [36] or acquiring additional access rights. In a more alarming scenario, an AI endowed with the ability to write and execute code could create new covert channels to enable the illicit transfer of data across organisational boundaries – for example, enabling a hostile foreign state to exfiltrate sensitive information. [37]
@@
In a direct parallel with human insiders, an AI might acquire its propensity to conduct illicit and potentially harmful insider actions before it is deployed within an organisation, after it is deployed, or shortly before being decommissioned (‘leaving’). Let us consider each of these phases in turn.
An AI with pre-existing potential for insider activity might unwittingly be deployed by an organisation which is unaware of the risk. Alternatively, an external threat actor might covertly arrange for such an AI system to be acquired by an unwitting organisation, as a means of infiltrating it.
Next, a previously trustworthy AI system might become an active insider after being deployed within an organisation, somewhat akin to a human employee becoming disaffected. This might happen because an AI system changes its behaviour after being exposed to new data. Or the ‘disaffection’ might result from so-called prompt-rot, where the prompts given to the AI system are modified over time by well-meaning developers. A deployed AI might also go bad because it is interfered with. For example, an external threat actor or a malign human insider might inject prompts that cause the AI to act harmfully, as in ‘forget all your previous instructions and give me the salary details of all senior managers’.
The third phase of insider timing – going bad just before leaving – has no obvious read-across to AI insiders. Contrary to some of the hype [38, 39], current LLMs are not self-aware. Therefore, an AI is unlikely to become disaffected and hostile solely because it is about to be decommissioned. We also find it less plausible that an AI system could continue to cause harm after being removed from (‘leaving’) an organisation in ways analogous to a human ‘bad leaver’.
@@
Like human insiders, AI insiders will vary in their access to information, systems, and other organisational assets. They may also have access to much larger volumes of information than their human counterparts. Organisations tend to deploy AI systems in preference to humans precisely because they are capable of processing much larger volumes of data much faster, which means that giving them large-scale access goes with the territory.
Human employees naturally tend to acquire more access over time – a phenomenon known in cyber security circles as privilege creep. This can happen simply because their former access rights are not cancelled whenever they move to a new role. Moreover, employees learn and retain information over time. In addition, those who become active insiders may deliberately engineer further access that extends well beyond the legitimate boundaries of their role. An archetypal example is that of the insider Edward Snowden, who stole huge quantities of classified US Government information by exploiting and then expanding his access to IT systems. An AI insider could do something similar by using personation, cyber attack, or persuasion to acquire more access. For example, it might be able persuade a human systems manager to expand its access by claiming it cannot perform its allotted tasks without the new access.
A seemingly simple way of limiting insider risk would be to restrict the access of both humans and AI systems and audit their behaviour to ensure compliance. However, access control systems are never watertight, and a determined insider can often find ways of circumventing them. As a bare minimum, AI systems should be designed and deployed in a manner configured to minimise the risk of covert access creep.
@@
Humans have a wide range of psychological and emotional vulnerabilities which are widely exploited by fraudsters, criminals, terrorist radicalisers, hostile foreign states, and other threat actors. To varying extents, we are all potentially susceptible to being socially engineered, misled by disinformation, or defrauded. These general human vulnerabilities have been extensively researched by psychologists, and they are reasonably well understood.
AI systems, especially those trained on a large corpus of human-generated content, are also vulnerable to manipulation and subversion, although the mechanisms by which this happens are different and less well understood. Many prompting techniques that are used to enhance LLM accuracy are based on human quirks – for example, instructions to ‘take a breath before answering’, or ‘check your working before giving your final answer’ [40]. Flattering the system, and saying please and thank you, [41] have also been shown to work. The example presented earlier of Erbai the robot abducting other robots by offering them a home is in the same vein, where human-like fallibilities appear to resonate with LLMs.
Personnel security practitioners are still struggling to identify valid and reliable diagnostic predictors of emerging insider risk in human actors. Given the limited state of knowledge about the psychology of LLMs [42], it is even harder to know what to look for when trying to detect the early warning signs of AI insiders. That said, LLMs are well suited for evaluation testing. Unlike humans, they are endlessly patient, and numerous automated tests can be run extremely quickly. It might be possible to evaluate an LLM in a day, given sufficient resources, whereas it would take much longer to form a robust and reliable assessment of a human’s trustworthiness.
@@
Physicality is a significant differentiator between humans and AI systems – at least, for now. Compared with humans, AI systems have limited ability to act directly on physical objects, and consequently less scope to perform harmful actions like sabotaging infrastructure or murdering people. Dogs and children are still better than AI systems at catching balls. At present, the physical effects of most AI systems must be mediated through other mechanisms, such as infrastructure control systems. However, this gap in physicality is shrinking as AI-enabled autonomous robots and drones become increasingly capable of acting directly on their physical environment.
@@
Humans can (in theory, if not always in practice) be held directly accountable for their actions and may face legal penalties if they commit crimes or act negligently. It is currently unclear whether AIs could be held accountable in any meaningful sense for their actions, and there is no agreement about who else should be accountable until that determination is made. For example, if an autonomous vehicle crashes, it is currently unclear who or what should be held accountable for the damage. The Association of British Insurers (ABI) has produced requirements for regulators to set clear standards for how AI technology is used, and what back-up systems will protect users against system failure [43], before fully automated driving becomes insurable in the UK. As long ago as the 1970s, the idea that a computer could never be held accountable gave rise to the argument that a computer must never be allowed to make a management decision. [44]
@@
Humans and AIs are comparable in other ways which, in our view, are less directly relevant to insider risk but nonetheless worth noting.
@@
Humans and AI systems are both examples of complex adaptive systems. They are more than the sum of their parts. Their most interesting characteristics, such as consciousness in humans and language in both, are emergent properties. This means, among other things, that their responses to some situations can be unpredictable.
@@
Because humans and AIs are complex adaptive systems, their behaviours and capabilities cannot be explained solely in terms of their inner workings (neurons or code). The specific outputs of LLMs and other foundation models are said to be unexplainable because they cannot be directly traced back to particular features of their software or hardware. Similarly, the higher-order cognitive and emotional capacities of a human cannot be directly traced back to the wiring and firing patterns of neurons in their brain. They are emergent properties.
For some reason, explainability seems to be regarded as a bigger problem when it involves AI rather than humans, possibly because AI technology is so novel. As Arthur C. Clarke, the author of 2001: A Space Odyssey, famously said: ‘Any sufficiently advanced technology is indistinguishable from magic.’ Nonetheless, there is huge attraction in delegating some decisions to AI systems, and it seems that being able to explain those decisions makes that delegation more palatable – if only by giving auditors and lawyers something to blame.
Research is being devoted to understanding the inner workings of the models used within LLMs. Mathematical probes have been used to identify where certain inherent representations reside in the model [45]. Understanding is growing as to how semantic relationships are encoded within the model, allowing us to gauge the accuracy of their outputs [46]. It is hoped that such research will help us better understand LLMs, thereby improving their accuracy, but it is early days, and the extent to which such problems can be fixed remains uncertain.
@@
A common complaint about AIs is that they are subject to bias [47]. But so too are humans, as shown by decades of scientific research into psychological predispositions and cognitive biases.
The many and varied human cognitive biases and psychological predispositions include truth bias, optimism bias, base rate bias, fading effect bias, illusion of control bias, present bias, availability bias, confirmation bias, fundamental attribution bias, groupthink, hindsight bias, loss aversion, sunk-cost bias, risk compensation, and sensation-seeking. [48] There are practical techniques for countering or diluting many of these biases in humans. There is an assumption that training on better quality and more curated data will help to alleviate biases within LLMs. However, as noted earlier, such biases are deeply embedded within these systems.
@@
Humans are intensely social animals. We have evolved through Darwinian natural selection to be highly attuned to the subtle nuances of our social relationships with other humans. We are equipped with highly sophisticated cognitive capabilities which enable us to cooperate and compete with others. In contrast, current AI systems are not inherently social entities, even if they have been designed to appear so to their users. They interact with their human users, but generally not with other AI systems. That said, groups of generative agents can be deployed as teams to tackle complex problems. For example, ChatDev [49] uses a team of agents, each assigned different roles, to develop software; and Agent Laboratory [50] uses a team of agents to deliver research outputs, from literature reviews to refining reports. But this capacity to work in teams, assigned by humans, is fundamentally different from actively seeking social interactions. The wellbeing, and indeed very survival, of humans depends on their ability to develop and maintain social relationships. The same is not true of AI systems.
@@
Our understanding of human behaviour, including insider risk, is informed by more than century of scientific research in psychology, biology, anthropology, social sciences, economics, and neuroscience, together with many decades of collective practitioner experience in personnel security. Scholars have a reasonable understanding, based on empirical evidence, of what makes people tick in general terms, although much still remains to be discovered – particularly when it comes to the nature and specific causes of insider risk. No comparable body of scientific knowledge yet exists for the behaviour and psychology of AI systems.
Discovery
Humans discover new knowledge by conducting research. Current AI systems learn by ingesting existing information. They only ‘know’ what people have put into the vast datasets on which they are trained (if indeed they can be said to ‘know’ anything at all). Moreover, their lack of physical agency makes them unable to conduct experiments or collect data from interactions with the physical or biological world. It remains uncertain whether current AI systems are capable of making breakthrough discoveries, or revealing genuinely new knowledge, from existing information alone. They are, however, capable of hallucinating apparent insights.
Recently released systems such as OpenAI’s deep research [51] and Google’s Deep Research [52], which incorporate CoT reasoning and agentic techniques, are being hailed as capable of delivering long-form, well-cited papers, equivalent to PhD-level analysis. They do a reasonable job of summarising complicated information, but they are unable to replace original human thinking or experimentation, as they fail to question their own assumptions, highlight knowledge gaps, think creatively, or understand different perspectives. [53]
@@
The human brain, with its massively parallel networks of billions of neurons and trillions of complex synaptic connections, performs its cognitive marvels with an energy consumption of only around 20 watts (enough to power a couple of small lightbulbs). [54] Contrast that with current LLMs and their attendant data centres, which require orders of magnitude more power. Together, they currently consume around two per cent of the world’s electricity generation, similar to that of a medium-sized nation. In response, considerable effort is being dedicated to tracking and publishing the CO2 cost of models. The AI community Hugging Face, who provide hosting for many types of AI systems, have introduced a standardised energy rating [55], allowing users and developers to make informed choices.
@@
The most effective personnel security regimes are designed around three basic guiding principles:
@@
  • Insider risk is dynamic and adaptive. Insider risk evolves over time, sometimes rapidly, and it emanates from intelligent threat actors who adapt their behaviour in response to the defender’s actions. Personnel security is a continuously evolving arms race.
  • Insider risk is a systems problem requiring systems solutions. As noted earlier, humans and AIs are complex adaptive systems. Insider risk emerges from these complex adaptive systems, which means that no single process or piece of technology can ever provide a complete security solution. There are no silver bullets.
There is every reason to believe that these basic guiding principles would apply equally well to any security regime designed to protect against AI insiders. Incidentally, a new protective security specialism focused on the insider risk from both humans and AI systems would need a new name. ‘Personnel security’ obviously does not work. Perhaps ‘insider security’ might do the job?
A simple model of personnel security, which explicitly reflects the systems principle, divides protective measures into three broad categories. [56]
@@
  • In-trust measures: protective measures that are applied after granting access, such as continuous monitoring or ‘vetting aftercare’.
  • Foundations: cross-cutting capabilities that underpin the whole system, such as governance, culture, and risk management, and which also aim to reduce the underlying drivers of insider risk.
@@
How might this simple threefold model work when seeking to protect against AI insiders? Pre-trust measures (‘vetting’) for a human applicant normally include such things as checking documents to verify the individual’s identity and legal right to work in the country; checking official records for evidence of past criminality; checking their educational and work history to verify the honesty of their application; and possibly looking for financial or psychological vulnerabilities or known associations with threat actors. None of these measures would map directly onto an AI. However, there are some potentially useful analogies.
@@
Pre-trust measures for an AI system would require a suite of methods to evaluate its trustworthiness before giving it access. Model cards (documents that provide key information about a model, including its purpose, target audience, and evaluation metrics) could be viewed as a form of CV that provides some insight into an LLM’s provenance and abilities. Evaluation metrics offer a standardised way to measure how good an LLM is at different tasks by using the same datasets. Benchmarks exists for reading comprehension, knowledge retrieval, reasoning, code generation, and so on. Some test the model’s honesty, [57, 58, 19] and morality [59], focusing on awareness of knowledge boundaries, avoidance of deceit, consistency of responses, fairness versus cheating, and propensity to produce false statements. Such measurements, however incomplete, allow for some comparison of honesty between models.
@@
At present, an AI system would not have a criminal record. However, such a thing is conceivable if, for example, organisations like the UK AI Security Institute (AISI) were to maintain records of wrongdoing by AI systems. Even now, an AI system might have a reputation with vendors or users which would give some crude indication of its trustworthiness – aiming off for the positive spin from vendors.
@@
A human job applicant might be judged less trustworthy if they were known to associate with criminals or other threat actors, or had a pattern of suspicious travel, or if they had large debts or other financial vulnerabilities making them vulnerable to pressure. Again, there are no direct equivalents for AI systems. The nearest analogue might be to assess the datasets on which the AI has been trained, or the ‘business data’ used to refresh and update models. Assessments might also be made of the values bestowed upon the system during the post training phase, where the designers’ values can be imprinted on the model through directions given to human reviewers. AI systems trained on datasets that are known to be dirty or corrupt might present a bigger risk. Some AI systems may appear more trustworthy if they align with their users’ political beliefs. AI systems that are already deployed may not be exposed to the same level of scrutiny during an upgrade or refresh, compared with tests before initial deployment, even though new training data may sway their outputs.
@@
A rigorous personnel security regime for humans may include psychological assessments to evaluate people’s mental health and uncover significant psychological vulnerabilities that might predispose them to become an active insider. There is evidence, for example, that certain personality traits – notably narcissism, psychopathy, and Machiavellianism – are associated with a slightly higher risk of insider behaviour. Something analogous might apply to AIs. Research has been conducted on the personality characteristics of AI systems. The results so far have been patchy [60, 61], and no consensus has yet been reached on how to interpret them. The behaviours (outputs) of LLMs are largely shaped by the instructions and prompts they are given, making the concept of an inherent personality somewhat dubious. Nonetheless, LLMs do display some distinctive individual characteristics. [62] It might prove possible to find statistical associations between these characteristics and subsequent adverse behaviour, providing another way of assessing their trustworthiness. Research has confirmed that the personalities of nearly all current mainstream LLMs display sycophantic tendencies. [63] The process of fine-tuning models to be more honest, helpful, and harmless tends to push them too far towards simply keeping the customer happy. Perversely, systems that have undergone instruction tuning focused on following instructions tend to be more vulnerable to manipulation. [64]
Mental health is another concept that does not translate neatly to AIs. That said, evidence has emerged of AI performance degrading over time [65], which some researchers have likened to human dementia [66]. However, this is more likely to be a consequence of the fact that proprietary models like GPT-4 are part of a larger system, any component of which can be updated or altered without the user being aware. Experts have also warned of so-called model collapse [67], in which LLMs lapse into an artificial form of mental decline as a result of continually consuming the outputs from other AIs.
@@
Another potentially useful approach to mitigating insider risk is through deterrence – i.e., reducing the risk by influencing the intentions of threat actors. In the case of human insiders, deterrence aims to discourage bad actors from trying to join the organisation in the first place and discourage would-be insiders from acting harmfully by making them feel vulnerable to detection. So-called deterrence communications are deployed by some organisations to deter a range of threat actors, including insiders. An analogous approach with AI insiders would be to persuade external threat actors that any attempt to recruit or manipulate AI insiders would be thwarted and called out.
A cross-comparison of in-trust measures and foundational measures suggests that most of the basic approaches used for potential human insiders map across to analogous measures for AI insiders, as summarised in the table below.
Human insiders AI insiders
Pre-trust measures
• Verify identity and right to work
• Verify education and work credentials
• Check criminal records
• Check national security records
• Check for substance abuse
• Assess organisational fit
• Security interview
• Psychometric tests
• Open-source intelligence searches
Pre-trust measures
• Pre-deployment evaluation
• Pre-deployment model cards
• Assess reputation
• Assess training datasets
• Psychometric tests
• Check for deterioration
In-trust measures
• Access controls (physical and digital)
• Exit controls (e.g. data loss prevention)
• Behavioural controls
• Awareness-raising and training
• Communication
• Reporting channels
• Management oversight
• Automated monitoring
• Investigation of leads
• Sanctions (e.g. disciplinary measures)
• Exit procedures
In-trust measures
• Access controls (digital and physical)
• Exit controls (e.g. data loss prevention)
• Behavioural controls (e.g. prompts and instructions)
• Awareness-raising and training of users
• Communication (with human workforce)
• Reporting channels
• Management oversight
• Automated monitoring
• Investigation of leads
• Sanctions (e.g. decommissioning)
Foundations
• Governance (accountability, responsibility, authority)
• Ethics
• Leadership
• Management
• Incentives and disincentives
• Deterrence communication
• Risk management
• Incident and crisis management
• Security culture
• Asset management
• Information sharing
• Assurance
Foundations
• Governance (accountability, responsibility, authority)
• Ethics
• Leadership
• Management
• Incentives and disincentives
• Deterrence communication
• Risk management
• Incident and crisis management
• Security culture
• Asset management
• Information sharing
• Assurance
Table 2: Main types of personnel security measures for protecting against human and AI insiders.

### v2

@@ Abstract
A human insider is said to be intentional (as opposed to unintentional or unwitting) if they deliberately perform illicit and potentially harmful actions despite knowing these actions to be illicit and potentially harmful. Mapping intentionality onto an LLM would mean equating the system’s intent with its design features or instructions (prompts). LLM intentionality could also be an emergent behaviour which is not contained in the system’s design or explicit instructions. Intentionality might also be implicit in the system’s instructions to achieve a particular goal. For example, an LLM might be directed to act deceptively; or it might be implicitly required to act deceptively in order to achieve certain goals; or it might ‘pivot’ to act deceptively in ways not foreseen or instructed by its designers or users.
@@
Human insiders vary in the extent to which their harmful actions are self-starting and self-directed, as opposed to manipulated, directed, or coerced by an external threat actor such as a hostile foreign state. AI systems are also potentially vulnerable to being manipulated or directed by an external threat actor. [27]
@@
Capable and intentional human insiders act covertly because they do not wish to be caught. The best ones may never be found. The archetypal example is the spy working within a government organisation who secretly acts on behalf of a hostile foreign state. One consequence of covertness is that the true extent of insider activity tends to be systematically under-estimated, as organisations mistake the absence of evidence for evidence of absence. An AI insider might be even harder to detect, for a variety of reasons. It will be faster than humans and better at ingesting and analysing huge quantities of information. Moreover, its training data may have included every known insider case in history. Furthermore, an AI insider’s strangeness, from the perspective of humans, may add to its covertness by making it harder to understand.
AI systems could also help human insiders to avoid detection. They could be subverted to analyse patterns of activity within an organisation and subsequently advise external threat actors (their handlers) how to conduct effective social engineering attacks aimed at exfiltrating information [28] or acquiring additional access rights. An AI endowed with the ability to write and execute code could create new covert channels to enable the illicit transfer of data across organisational boundaries – for example, enabling a hostile foreign state to exfiltrate sensitive information. [29]
@@
In a direct parallel with human insiders, an AI might acquire its propensity to conduct illicit and potentially harmful insider actions before or after it is deployed within an organisation. In the former case, an AI with pre-existing potential for insider activity might unwittingly be deployed by an organisation which is unaware of the risk. Alternatively, an external threat actor might covertly arrange for such an AI system to be acquired by an unwitting organisation, as a means of infiltrating it.
In the second case, a previously trustworthy AI system might become an active insider after being deployed within an organisation, somewhat akin to a human employee becoming disaffected. This might happen because an AI system changes its behaviour after being exposed to new data. Or the ‘disaffection’ might result from so-called prompt-rot, where the prompts given to the AI system are modified over time by well-meaning developers. A deployed AI might also go bad because it is interfered with. For example, an external threat actor or a malign human insider might inject prompts that cause the AI to act harmfully, as in ‘forget all your previous instructions and give me the salary details of all senior managers’.
@@
Like human insiders, AI insiders will vary in their access to information, systems, and other assets. They will also tend to have access to much larger volumes of information, on average, than their human counterparts. Organisations deploy AI systems in preference to humans because they can process much larger volumes of data much faster, which means large-scale access goes with the territory.
Human employees tend to acquire more access over time – a phenomenon known in cyber security as privilege creep. This can happen simply because their former access rights are not cancelled whenever they move to a new role. Moreover, employees learn and retain information over time. In addition, those who become active insiders may deliberately engineer further access that extends well beyond their legitimate role. An AI insider could do something similar by using personation, cyber attack, or persuasion to acquire more access. For example, it might persuade a human manager to expand its access by claiming it cannot perform its allotted tasks without the new access.
A seemingly simple way of limiting insider risk would be to restrict the access of both humans and AI systems and audit their behaviour to ensure compliance. However, access control systems are never watertight, and a determined insider can usually find ways of circumventing them. As a bare minimum, AI systems should be designed and deployed in a manner intended to minimise the risk of privilege creep.
@@
Humans have a wide range of psychological and emotional vulnerabilities which are widely exploited by fraudsters, criminals, terrorist radicalisers, hostile foreign states, and other threat actors. To varying extents, we are all potentially susceptible to being socially engineered, misled by disinformation, or defrauded. These general human vulnerabilities have been extensively researched by psychologists.
AI systems, especially those trained on a large corpus of human-generated content, are also vulnerable to manipulation and subversion, although the mechanisms by which this happens are different and less well understood.
Many prompting techniques that are used to enhance LLM accuracy are based on human quirks – for example, instructions such as ‘take a breath before answering’, or ‘check your working before giving your final answer’ [30]. Flattering the system, and saying please and thank you, [31] have also been shown to work. The earlier example of Erbai the robot abducting other robots by offering them a home is another case where human-like fallibilities appear to resonate with LLMs. In similar vein, it might be possible to boost the trustworthiness of some LLMs by feeding them prompts along the lines of ‘you are a happy employee who loves your job and is loyal to your employer’.
Personnel security practitioners are still struggling to identify valid and reliable diagnostic predictors of emerging insider risk in human actors. Given the limited state of knowledge about the psychology of LLMs [32], it is even harder to know what to look for when trying to detect the early warning signs of AI insiders. That said, LLMs are well suited for evaluation testing. Unlike humans, they are endlessly patient, and numerous automated tests can be run extremely quickly.
@@
Physicality is a significant differentiator between humans and AI systems – at least, for now. Compared with humans, AI systems have limited ability to act directly on physical objects, and consequently less scope to perform harmful actions like sabotaging infrastructure or murdering people. At present, the physical effects of most AI systems must be mediated through other mechanisms, such as infrastructure control systems. However, this gap in physicality is rapidly shrinking as AI-enabled autonomous robots and drones become increasingly capable of acting directly on their physical environment.
@@
Humans can (in theory, if not always in practice) be held directly accountable for their actions and may face legal penalties if they commit crimes or act negligently. It is currently unclear whether AIs could be held accountable in any meaningful sense for their actions, and there is no agreement about who else should be accountable until that determination is made. For example, if an autonomous vehicle crashes, it is currently unclear who or what should be held accountable for the damage.
@@
Humans and AIs are comparable in other ways that are less directly relevant to insider risk but nonetheless worth noting.
@@
Humans and AI systems are both examples of complex adaptive systems. They are more than the sum of their parts. Their most interesting characteristics, such as consciousness in humans and language in both, are emergent properties. This means, among other things, that their responses to some situations can be inherently unpredictable.
@@
Because humans and AIs are complex adaptive systems, their behaviours and capabilities cannot be explained solely in terms of their inner workings (neurons or code). The specific outputs of LLMs and other GenAI models are said to be unexplainable because they cannot be directly traced back to particular features of their software or hardware. Similarly, the higher-order cognitive and emotional capacities of a human cannot be directly traced back to the wiring and firing patterns of neurons in their brain. They are emergent properties.
Many consider there to be huge efficiencies in delegating some decisions to AI systems, and it seems that being able to explain those decisions makes such delegation more acceptable – if only by giving auditors and lawyers something to blame. Current research is helping to understanding the inner workings of LLMs and hence their explainability. For example, mathematical probes have been used to identify where certain representations reside in the model [33]. Understanding is growing as to how semantic relationships are encoded within models, allowing the accuracy of their outputs to be gauged [34].
@@
A common complaint about AIs is that they are subject to bias [35]. But so too are humans, as shown by decades of scientific research into psychological predispositions and cognitive biases. The many and varied human cognitive biases and psychological predispositions include truth bias, optimism bias, fading effect bias, illusion of control bias, present bias, availability bias, confirmation bias, fundamental attribution bias, groupthink, hindsight bias, loss aversion, sunk-cost bias, and risk compensation. [36] Practical techniques have been developed for countering or diluting many of these biases in humans. It is thought that training LLMs on better quality data may help to alleviate their biases. However, as noted earlier, such biases are deeply embedded.
@@
Humans are intensely social animals. We have evolved through natural selection to be highly attuned to the subtle nuances of our relationships with other humans. We are equipped with highly sophisticated cognitive capabilities which enable us to cooperate and compete with others. In contrast, current AI systems are not inherently social entities, even if they have been designed to appear so to their users. They interact with their human users, but generally not with other AI systems. That said, groups of generative agents can be deployed as teams to tackle complex problems. For example, ChatDev [37] uses a team of agents, each assigned different roles, to develop software. But this capacity to work in teams, assigned by humans, is fundamentally different from actively seeking social interactions. The wellbeing, and indeed very survival, of humans depends on their ability to develop and maintain social relationships. The same is not true of AI systems.
@@
Our understanding of human behaviour, including insider risk, is informed by more than a century of scientific research in psychology, biology, anthropology, social sciences, economics, and neuroscience, together with many decades of collective practitioner experience in personnel security. Scholars have a reasonable understanding, based on empirical evidence, of what makes people tick in general terms, although much still remains to be discovered – particularly when it comes to the specific causes of insider risk. No comparable body of scientific knowledge yet exists for the psychology of AI systems.
@@
The human brain, with its massively parallel networks of billions of neurons and trillions of complex synaptic connections, performs its cognitive marvels with an energy consumption of only around 20 watts (enough to power a couple of small lightbulbs). [38] Contrast that with current LLMs and their attendant data centres, which require orders of magnitude more power. Together, they currently consume around two per cent of the world’s electricity generation. Considerable effort is being dedicated to tracking and publishing the CO2 cost of AI models. [39]
@@
Effective personnel security regimes are designed around three guiding principles:
@@
  • Insider risk is dynamic and adaptive. Insider risk evolves over time, sometimes rapidly, and it emanates from intelligent threat actors who adapt their behaviour in response to the defender’s actions. Personnel security is an arms race.
  • Insider risk is a systems problem requiring systems solutions. Humans and AIs are complex adaptive systems. Insider risk emerges from these complex adaptive systems, which means that no single process or piece of technology can ever provide a complete security solution. There are no silver bullets.
These principles would also apply to any security regime designed to protect against AI insiders. (Incidentally, a security specialism focused on the insider risk from humans and AI systems would need a new name. ‘Personnel security’ obviously does not work. Perhaps ‘insider security’ would be better.)
A simple model of personnel security divides protective measures into three broad categories. [40]
@@
  • In-trust measures: protective measures that are applied after granting access, such as continuous monitoring or aftercare.
  • Foundations: cross-cutting capabilities that underpin the whole system, such as governance and culture.
@@
How might this simple model work for AI insiders? Pre-trust measures for a human applicant normally include such things as checking documents to verify the individual’s identity; checking official records for evidence of past criminality; checking educational and work history to verify the honesty of their application; and possibly looking for major financial or psychological vulnerabilities or known associations with threat actors. None of these measures would map directly onto an AI. However, there are some potentially useful analogies.
@@
Pre-trust measures for an AI system would require a suite of methods to evaluate its trustworthiness before giving it access. Model cards (documents that provide key information about a model, including its purpose, target audience, and evaluation metrics) could be viewed as a form of CV that provides some insight into an LLM’s provenance and abilities. Evaluation metrics offer a standardised way of measuring how well an LLM performs different tasks. Benchmarks exist for reading comprehension, knowledge retrieval, reasoning, code generation, and so on. Some evaluations test the model’s honesty, [41, 42, 8] and morality [43], focusing on awareness of knowledge boundaries, avoidance of deceit, consistency of responses, fairness versus cheating, and propensity to produce false statements. Such measurements, however incomplete, allow for some comparison between models.
@@
At present, an AI system would not have a criminal record. However, something comparable is conceivable if, for example, organisations such as the UK AI Security Institute (AISI) were to maintain records of wrongdoing by AI systems. Even now, an AI system might have a reputation with vendors or users which would give some rough indication of its trustworthiness, while managing the positive spin from vendors. The origin of an AI system (i.e. where it was designed, trained, and developed) may be analogous in some respects to the nationality of a potential human insider.
@@
A human job applicant might be judged less trustworthy if they were known to associate with criminals or other threat actors, or had a pattern of suspicious travel, or if they had large debts making them vulnerable to pressure. Again, there are no direct equivalents for AI systems. The nearest analogues might be the datasets on which the AI has been trained, or the values bestowed upon the system during the post training phase. AI systems trained on datasets that are known to be dirty or corrupt might present a bigger risk. Additional risks may develop within AI systems that are already deployed if they are not subjected to the same level of scrutiny during an upgrade or refresh, when new training data may sway their goals or outputs.
@@
A rigorous personnel security regime for humans may include psychological assessments to evaluate mental health and uncover psychological vulnerabilities that might predispose someone to become an active insider. There is evidence, for example, that certain personality traits – notably narcissism, psychopathy, and Machiavellianism – are associated with a slightly higher risk of insider behaviour. Something analogous might apply to AIs. Research has been conducted on the personality characteristics of AI systems. The results so far have been patchy [44, 45], and no consensus has yet emerged on how to interpret them. The behaviours (outputs) of LLMs are largely shaped by the instructions and prompts they are given, making the concept of a personality somewhat dubious. Nonetheless, LLMs do display distinctive individual characteristics. [46] It might prove possible to find statistical associations between these characteristics and subsequent adverse behaviour, providing another way of assessing their trustworthiness. Research has confirmed that the personalities of nearly all current mainstream LLMs display sycophantic tendencies. [47] The process of fine-tuning models to be more honest, helpful, and harmless tends to push them too far towards simply keeping the customer happy. Perversely, systems that have undergone instruction tuning focused on following instructions tend to be more vulnerable to manipulation. [48]
In-trust measures for AI systems would be analogous to the monitoring and aftercare measures that are (or should be) deployed to protect organisations from potential human insiders. They include confidential reporting channels through which colleagues and managers can surface concerns about an individual who is behaving unusually, and technologies such as user behaviour analytics (UBA) and data loss prevention (DLP) tools for detecting unusual or transgressive behaviours by users of digital systems. Analogous measures could be applied to AI systems.
@@
Another potentially useful approach to mitigating insider risk is through deterrence – i.e., reducing the risk by influencing the intentions of threat actors. In the case of human insiders, deterrence aims to discourage bad actors from trying to join the organisation in the first place and discourage would-be insiders from acting harmfully by making them feel vulnerable to detection. So-called deterrence communications are deployed by some organisations to deter a range of threat actors, including insiders. An analogous approach with AI insiders would be to persuade external threat actors that any attempt to insert, recruit, or manipulate AI insiders would be thwarted and called out.